Sample text for slashbox.sty

The usage is pretty straightforward, such as

[figure: Room; Date]

5/31 6/1 6/2 6/3

6/4

Meeting Room Auditorium Seminar Room

You may include a newline (\\) in ‘Room’ and/or ‘Date’. Note that you will get spaces aside the slash line if there is a wider column in the same column of a different line. In such a case, you need to specify the width of the slashed column by saying

[figure: Room; Date]

5/31 6/1 6/2

6/3

Room With a Long Name Auditorium Seminar Room

The specified width is neglected if it is narrower than the natural width of the column.

\(back)slashbox assumes by default that there is a blank space of width \tabcolsep on both sides of the column. Thus the slash line might exceeds the boundary when you use @{} etc.

You can avoid it by specifying

[figure: Room; Date]

5/31 6/1 6/2 6/3

6/4

• Meeting Room

• Auditorium

• Seminar Room

Here [l] tells the command that there is no extra space on the left of this column. You can use [r] and [lr] likewise. You have to also specify the width of the column in this case, but it can be 0pt.